Hi, Nick again.
I have another question, this time regarding how to "Make t the subject of the expresion"- please see below equation(s).
Equation 1
A/(squareroot(t-b)) =c
Equation 2
h= ln(M/((1/3t^2)+f))
-How do I make t the subject of the above equations. Steps and explanations as-well as solutions would be most appreciated.
Thanks, -Nick.

Answer by stanbon(75887)

"Make t the subject of the expresion"- please see below equation(s).
Equation 1
A/(squareroot(t-b)) =c
---
sqrt(t-b) = c/A
---
t-b = (c/A)^2
t = (c/A)^2 + b
=================================
Equation 2
h= ln(M/((1/3t^2)+f))
h = ln(M) - ln(((1/3t^2)+f))
ln((1/3t^2)+f)) = ln(M) - h
---
(1/3)t^2+f = e^(ln(M) - h)
---
t^2 = 3[e^(ln(M) - h)]-f
---
t = +/-sqrt[3[e^(ln(M) - h)]-f]
=====================================
Cheers,
Stan H.